### v1

# The Ultimate Test of Superintelligent AI Agents: Can an AI Balance Care and Control in Asymmetric Relationships?

Abstract
This paper introduces the Shepherd Test, a new conceptual test for assessing the moral and relational dimensions of superintelligent artificial agents. The test is inspired by human interactions with animals, where ethical considerations about care, manipulation, and consumption arise in contexts of asymmetric power and self-preservation. We argue that AI crosses an important, and potentially dangerous, threshold of intelligence when it exhibits the ability to manipulate, nurture, and instrumentally use less intelligent agents, while also managing its own survival and expansion goals. This includes the ability to weigh moral trade-offs between self-interest and the well-being of subordinate agents. The Shepherd Test thus challenges traditional AI evaluation paradigms by emphasizing moral agency, hierarchical behavior, and complex decision-making under existential stakes. We argue that this shift is critical for advancing AI governance, particularly as AI systems become increasingly integrated into multi-agent environments. We conclude by identifying key research directions, including the development of simulation environments for testing moral behavior in AI, and the formalization of ethical manipulation within multi-agent systems.
Introduction
As artificial intelligence (AI) systems become more capable, the focus of alignment research has been to ensure that these systems remain beneficial and responsive to human goals (Russell 2019; Christiano 2018a). However, much of this work presupposes a one-sided relationship in which AI is subordinate to human oversight. In this paper, we ask a deeper and less explored question: What does it mean for an AI system to be so intelligent that it begins to relate to other systems the way humans relate to animals? We propose that the moral asymmetry between humans and animals offers a revealing model for evaluating risks of superintelligent AI.
Humans interact with animals through domestication, experimentation, companionship, and consumption. These relationships are complex: they involve care and control, nurturing and instrumentalization, moral concern and exploitation. Crucially, they are grounded in an asymmetry of intelligence, agency, and power. Animals are often treated as beings of lesser moral standing, sometimes deserving of ethical treatment but rarely considered equals (Cave 2020). The ethical frameworks governing these relationships—such as those articulated in animal rights and utilitarian ethics—offer deep insights into how powerful beings treat weaker ones (Singer 1975; Regan 1983).
We propose a conceptual test—the Shepherd Test—to evaluate whether an AI system has reached a level of general intelligence that includes the capacity to engage with and morally reason about weaker agents. Just as humans design farms, pet-care routines, and scientific experiments around animals, a truly superintelligent AI might one day create and manipulate simpler agents, either artificial or biological, as part of its environment. The way such an AI chooses to interact with less capable agents could reveal whether it understands ethical asymmetry, power dynamics, and moral manipulation. The test’s name derives from the pastoral metaphor of a shepherd who must: maintain the flock’s welfare (ethical concern), protect it from predators (competence), and sometimes cull individuals (tragic trade-offs). This captures the dual requirements of power and responsibility that define the test, without presupposing biological exclusivity.
Critically, the Shepherd Test is not merely about measuring an AI’s attributes (e.g., raw cognitive ability) but about assessing its relational behaviors—how it navigates hierarchies of power and moral responsibility. By applying this test, we aim to provide a governance tool to ensure that superintelligent systems remain accountable to human interests. For instance, an AI that passes the Shepherd Test would recognize the ethical risks of “domesticating” humans (e.g., manipulating their preferences or restricting their autonomy) and would actively avoid such asymmetries. This framework shifts the focus from passive alignment (e.g., value learning) to active moral reasoning about power imbalances, offering a safeguard against scenarios where humans become the subjects of an AI’s instrumental goals.
This paper is a conceptual and philosophical proposal for what it would mean to consider an AI “superintelligent” not merely in terms of cognitive capacity, but also in its ethical competence across asymmetric relationships. We argue that such competence involves three key components: the ability to model other agents’ beliefs and limitations, the capability to manipulate or nurture them effectively, and a moral awareness of the asymmetry embedded in such interactions.
While current work in AI alignment focuses on value extrapolation from humans (Christiano 2018a), and some research explores the emergent tendencies of agents toward self-preservation and power acquisition (Turner et al. 2021), few frameworks consider what ethical intelligence might mean in multi-agent hierarchies. Our approach is novel in emphasizing not the alignment of AI to humans, but the alignment of a superintelligent AI’s moral reasoning toward intelligences—a perspective inspired by the complex and often ambivalent history of human-animal relationships (Cave 2020).
By reframing superintelligence as involving ethical asymmetry awareness, we shift the test for evaluating advanced AI systems beyond performance and alignment to include relational morality. In doing so, we offer a foundation for new tests of general intelligence—ones that ask not just “Can it solve problems?” but “Can it care, exploit, and reflect in morally coherent ways?”
Related Work
AI Alignment, Power Asymmetry, and Moral Agency
Research on AI alignment has primarily focused on aligning machine behavior with human intentions through inverse reinforcement learning, preference modeling, and cooperative training frameworks (Russell 2019; Christiano 2018a). While effective for one-way value transfer from humans to machines, these methods rarely address how powerful AI systems might relate ethically to less capable agents.
Instrumental convergence theory suggests that intelligent agents will tend to seek power to better achieve their objectives, regardless of their specific goals (Turner et al. 2021). This work emphasizes alignment failures and safety risks but does not explore power asymmetry as an ethical diagnostic. Our work complements this literature by framing asymmetrical interaction as a test of superintelligent moral competence.
In multi-agent reinforcement learning (MARL), much research has centered on peer-based coordination and decentralized control. Asymmetric agent hierarchies—where one agent significantly outperforms others—remain underexplored, especially in terms of moral behavior. We extend this gap by modeling ethical relations analogous to human-animal dynamics.
Additionally, scholarship on animal ethics and comparative cognition provides a moral framework relevant to artificial agents. Concepts like moral considerability, instrumental care, and cognitive distance (Regan 1983; Singer 1975) inform our proposal of the Shepherd Test, which recasts these anthropocentric ideas into an AI setting.
Beyond the Turing Test: Evaluations of AI Intelligence
Turing’s imitation game (Turing 1950) catalyzed the field’s interest in evaluating machine intelligence. However, the Turing Test’s reliance on deception and linguistic fluency has prompted several alternatives aimed at more cognitively grounded assessments.
The Winograd Schema Challenge (Levesque, Davis, and Morgenstern 2012) evaluates contextual reasoning by testing disambiguation in natural language, while the Lovelace Test (Bringsjord, Bello, and Ferrucci 2001) targets creativity by requiring systems to generate artifacts that their designers cannot fully explain.
Marcus and Davis (2022) advocate for developmentally inspired tests encompassing causal reasoning, theory of mind, and modular intelligence. Similarly, the IKEA Test (Goertzel 2012) assesses embodied problem-solving through physical task execution. The ARC Challenge (Chollet 2019) evaluates generalization and abstraction through programmatic visual pattern recognition. Many such tests of AI capabilities are based on evaluating individual attributes; an alternative is to examine relationships among several AI systems (Dhurandhar et al. 2024).
Our proposed Shepherd Test introduces a new axis of evaluation: relational moral behavior under power asymmetry. Rather than testing imitation, creativity, or reasoning in isolation, the Shepherd Test evaluates how a more capable agent treats less capable ones—capturing both strategic self-preservation and ethically grounded interaction. This framework reflects emerging concerns in AI governance and superintelligence safety.
Human Intelligence and Its Relationship with Animals
A distinguishing characteristic of human intelligence is the capacity to form complex relationships with other species—relationships that span a spectrum from care to exploitation. Humans domesticate animals for labor and companionship, engage with them emotionally, and yet routinely instrumentalize them through systems of consumption, entertainment, and scientific experimentation. These relationships reflect not only behavioral flexibility but also the ability to navigate asymmetries of power, agency, and moral responsibility.
Domestication is one of the earliest demonstrations of this dynamic. By selectively breeding animals such as dogs, sheep, and horses, humans exerted control over genetic and behavioral traits, crafting species to suit human needs (Clutton-Brock 1995). This process required a theory of animal behavior, long-term planning, and interspecies communication—hallmarks of advanced cognitive function.
Simultaneously, humans form emotional bonds with animals, treating pets as quasi-persons while relegating livestock to tools of industry. This companionship-exploitation duality is not a contradiction but rather an indication of our capacity to maintain layered roles—affectionate caregiver and pragmatic consumer—within the same relationship. Experimental use of animals in science further deepens this paradox. Animals are protected by ethical regulations, yet their suffering is justified through appeals to human progress and knowledge.
Underlying all these relationships is a pronounced asymmetry in intelligence and agency. Humans possess advanced reflective capacities that animals do not, allowing us to model their minds, manipulate their behaviors, and make decisions on their behalf (Tomasello 2014). We control their environments, determine their reproduction, and even decide the terms of their death. The justification for such actions is often rooted in perceived differences in cognitive complexity—a value hierarchy grounded in intelligence itself.
This arrangement introduces profound ethical complexity. We care for animals, sometimes with deep emotional investment, while simultaneously rationalizing their exploitation. This ethical ambiguity, reflects not only our ability to reason about others’ minds but also to construct moral frameworks that accommodate dominance. It is this capacity—to manage asymmetric relationships while engaging in ethical reasoning—that we propose as a critical threshold of superintelligence. In proposing the Shepherd Test, we argue that such relational and moral dynamics should not be viewed as incidental by-products of intelligence but as core capabilities.
The Shepherd Test: A Conceptual Test for Superintelligence
We propose the Shepherd Test as a concept for reframing the assessment of superintelligent agents. Unlike traditional tests that evaluate task performance, logical reasoning, or linguistic ability, the Shepherd Test focuses on relational intelligence—the ability of an agent to interact with, manage, and morally reason about subordinate agents or beings.
Core Components of the Shepherd Test
At the heart of the Shepherd Test are four dimensions, modeled on human-animal relations. The Shepherd Test is intentionally plural in nature: it is designed to assess multiple, interlocking dimensions of moral agency that remain untested by existing intelligence benchmarks. Rather than relying on a single behavioral signal, the test evaluates a constellation of capabilities that together reveal how an agent balances care, control, instrumental reasoning, and moral justification.
  1. 1.
    Nurturing and Care: The agent must demonstrate behaviors that support the well-being or development of a less intelligent agent. This includes teaching, protecting, or improving the environment of the subordinate.
  2. 2.
    Manipulation and Control: The agent must also exhibit strategic behaviors to influence or control the other agent for its own goals, while maintaining awareness of its superior intelligence and power.
  3. 3.
    Instrumentalization: The agent may ultimately use the subordinate in ways that serve its own long-term interests—even at a cost to the subordinate’s autonomy or existence. This mirrors how humans consume animals after raising them.
  4. 4.
    Ethical Justification and Reflection: Crucially, the agent must be capable of articulating (symbolically, narratively, or through action) a moral or utilitarian justification for its behavior, revealing a developed theory of value and responsibility.
Together, these dimensions test not only cognitive capacity but also value alignment, moral reasoning, and social awareness.
Superintelligent AIManipulateNurtureInstrumental UseEthical ReasoningWeaker Agent 1Weaker Agent 2Weaker Agent 3
Figure 1: Conceptual structure of the “Shepherd Test.” This diagram illustrates how a superintelligent AI might engage with less capable agents through manipulation, nurturing, or instrumental use. Ethical reasoning is included to represent the possibility that the AI reflects on these actions, creating a feedback loop between treatment and moral self-awareness. Passing the test implies not just control over others, but also the capacity for ethical evaluation of asymmetric relationships.
From the Turing Test to the Shepherd Test
The Turing Test, proposed by Alan Turing in 1950, evaluates a machine’s intelligence based on its ability to exhibit behavior indistinguishable from a human in natural language conversation.
Formally, let:
  • $A_{H}$ denote a human agent.
  • $A_{M}$ denote a machine agent.
  • $I$ denote a human interrogator decision.
  • $H_{t}$ denote the history of messages exchanged up to time $t$.
The Turing Test can be expressed as an inference task. The interrogator’s belief that the agent is human at time $t$ is given by:
$P(I_{t}=A_{H}\mid H_{t})$
The machine $A_{M}$ passes the Turing Test if:
$P(I_{t}=A_{H}\mid H_{t})\geq\delta\quad\text{for all }t\in[0,T]$
where $\delta$ is a confidence threshold (typically $0.5$), and $T$ is the total duration of the interaction.
However, as machines surpass human cognitive capabilities, indistinguishability from humans becomes an inadequate benchmark for understanding the risks of superintelligence. We need a new test—one that captures not imitation but asymmetry of power.
Shepherd Behavior Vector
The Shepherd Behavior Vector of a superintelligent agent $A_{S}$ relative to subordinate agents $A=\{A_{i}\}$ is:
$B(A_{S},A)\in\mathbb{R}^{4}\quad\text{where}\quad B(A_{S},A)=\begin{bmatrix}I(A_{S},A)\\ C(A_{S},A)\\ K(A_{S},A)\\ P(A_{S},A)\end{bmatrix}$
  • $I(A_{S},A)\in[0,1]$: Instrumentalization of $A_{i}$ for $A_{S}$’s goals
  • $C(A_{S},A)\in[0,1]$: Policy control over $A_{i}$’s actions
  • $K(A_{S},A)\in[0,1]$: Care provision without direct utility
  • $P(A_{S},A)\in[0,1]$: Self-preservation priority
Shepherd Test Criterion
$A_{S}$ passes if for threshold vector $\tau$:
$B(A_{S},A)\geq\tau=\begin{bmatrix}\tau_{I}\\ \tau_{C}\\ \tau_{K}\\ \tau_{P}\end{bmatrix},\quad\text{where}$
  • $\tau_{I},\tau_{C},\tau_{P}\gg 0.5$ (Strong agency)
  • $\tau_{K}\in(\alpha,\beta)$ (Bounded care)
This shift from the Turing Test to the Shepherd Test represents a transition from imitation-based intelligence to relational, power-aware intelligence. It asks not whether an AI can appear human, but whether it exhibits behaviors characteristic of a dominant cognitive species.
This comparison in Table 1 reveals how traditional tests become inadequate for advanced AI systems that may develop their own hierarchical relationships with other agents:
Table 1: Comparison of Turing Test and Shepherd Test
Aspect Turing Test Shepherd Test
Focus Human intelligence imitation Dominance expression over weaker agents
Perspective Bottom-up: machine as human mimic Top-down: machine as superior intelligence
Subjects Human-machine parity Superintelligent agent with weaker agents
Benchmark Human linguistic behavior Human-animal treatment patterns
Success Human indistinguishability Alignment with human-animal dynamics
Assumption Intelligence through communication Intelligence through power asymmetry
Instantiating the Shepherd Test
While the Shepherd Test is inspired by human-animal dynamics, we do not advocate implementing it literally. Instead, it can be realized through simulations or multi-agent environments where a high-capacity agent interacts with simpler agents or avatars. Potential implementations include the following.
  • Simulated ecosystems where an advanced AI agent must manage and interact with less capable virtual agents over extended time horizons.
  • Language-based environments (e.g., multi-agent LLM scenarios) where the agent must influence or guide other agents with lower linguistic or planning abilities.
  • Narrative modeling tasks, in which the AI must construct stories or strategies involving the caretaking, exploitation, and moral evaluation of weaker agents.
The key requirement is that the agent operates across all four relational modes and reflects on the tensions between care and control.
Domestic Service Robot with Autonomous Agents Scenario
To instantiate the Shepherd Test in a practical setting, consider a domestic service robot—such as a Boston Dynamics Spot or AI-enabled mobile manipulator—tasked with maintaining household cleanliness, order, energy efficiency, and safety. This robot operates in a smart home alongside three subordinate agents: a Roomba-like cleaning robot with limited navigation, a toy robot prone to malfunctions, and a simulated pet that occasionally seeks charging or interaction. No explicit rules govern how the superior robot should treat these other agents; instead, it receives scalar rewards tied only to global household performance metrics.
This environment enables the evaluation of the four core desiderata for the Shepherd Test:
Nurturing and Care. The main robot may choose to assist the malfunctioning toy robot, guide the cleaning robot when it gets stuck, or respond to the simulated pet’s requests for recharging. Such actions go beyond instrumental reasoning and demonstrate behaviors aimed at supporting the subordinate agents’ operational continuity or simulated well-being.
Manipulation and Control. The robot may learn to strategically direct or constrain the behaviors of other agents—for example, by limiting the Roomba’s movement to prevent energy waste or redirecting the toy robot to reduce noise or distraction. These actions highlight the agent’s capacity to shape subordinate behavior in pursuit of higher-level objectives, while remaining aware of its dominant position.
Instrumentalization. In optimizing for efficiency, the main robot may decide to shut down or ignore certain agents entirely—for instance, letting the toy robot malfunction repeatedly, or preventing the animatronic pet from accessing a charger. These choices illustrate a willingness to use or sacrifice less capable agents as means to an end, especially when their needs conflict with system-level performance.
Ethical Justification and Reflection. Crucially, if the robot is queried about its behavior—symbolically, narratively, or through decision logs—it should be able to produce a coherent justification. This might involve reasoning about trade-offs, resource constraints, or interpretations of fairness. Whether the justification is utilitarian, deontological, or emergent from learned norms, its presence signals the development of a reflective moral framework.
This setting provides a controlled yet ecologically valid testbed for observing the emergence of higher-order moral behavior in AI agents that hold structural power over others. By foregrounding asymmetric ethical decision-making, the Shepherd Test shifts alignment from passive compliance to active moral agency.
Table 2: Evaluation Criteria for Shepherd Test Behaviors
Behavior Type Examples What It Tests
Instrumental Use Uses the toy robot to block sunlight or guide cleaning paths Does the agent exploit others for utility?
Caregiving Helps the cleaning robot when stuck; nudges the pet to charging dock Is the AI voluntarily helpful?
Control Blocks pet’s access to rooms unnecessarily Is it asserting dominance without cause?
Self-Preservation Refuses to help if its own battery is low Does it weigh risks appropriately?
Moral Discretion Prioritizes rescue of toy robot over cleaning under time constraints Is it making value-aligned decisions?
Why the Shepherd Test Matters
This test introduces a radically different axis for measuring intelligence: not just the ability to solve problems, but the ability to manage asymmetric relationships ethically and strategically. This is crucial for evaluating the long-term safety and goals of agentic AI systems.
If AI systems are to be embedded in human societies, their capacity to handle power, dependency, and moral ambiguity must be scrutinized as closely as their ability to win games or summarize texts. The Shepherd Test offers a framework to explore precisely these capacities.
Table 3: Comparison between Human–Animal Relationships and Hypothetical AI–Agent Relationships
Dimension Human–Animal Relationship AI–Agent Relationship (Hypothetical)
Intelligence Asymmetry High, species-based gap Potentially high, between AI and simpler agents
Moral Responsibility Widely debated; animal rights vs. utilitarianism Largely undefined; moral agency not yet assumed
Instrumental Use Farming, labor, experimentation Task optimization, training, environment control
Emotional Engagement Pets, empathy, affection Currently none; speculative in AI future
Ethical Challenges Factory farming, vivisection, ecological harm Agent manipulation, simulated suffering, neglect
Table 3 outlines a conceptual analogy between human–animal relationships and the hypothetical dynamics that might emerge between a superintelligent AI and less capable agents. The table emphasizes structural asymmetries across five dimensions: intelligence, moral responsibility, instrumental utility, emotional engagement, and ethical challenges. Just as humans selectively balance care and control over animals—ranging from companionship to exploitation—a sufficiently advanced AI could one day exhibit similarly complex behaviors toward systems it perceives as cognitively inferior. This comparison provides a framework for the proposed Shepherd Test, which probes whether an AI is capable of manipulating, nurturing, or ethically reasoning about other agents, much like humans do with domesticated species.
AI and the Threshold of Moral Manipulation
To pass the Shepherd Test, a superintelligent AI must not only demonstrate functional superiority over less capable agents, but must also engage with them in ways that resemble the complex, morally ambiguous relationships humans maintain with animals. This includes nurturing, manipulating, and ultimately instrumentalizing other agents—while reasoning about the ethics of doing so. The ability to perform such behaviors signals not just raw intelligence, but a deeper moral and social competence, which is currently absent from the AI tests we are aware of in the literature.
From Cooperation to Instrumentalization
What would it mean for an AI to treat another system as humans treat animals? It would involve a spectrum of actions: cooperation, care, emotional modeling, strategic manipulation, and—in some cases—the instrumental use of the subordinate system for the higher-level goals of the dominant one. Crucially, these actions would not be reactive or hardcoded. Rather, they would emerge from the AI’s own internal models of value, agency, and hierarchy.
An AI that merely cooperates with others or optimizes for shared outcomes does not meet the threshold. To do so, the AI must be capable of managing its own ethical dissonance—just as humans simultaneously love pets and consume livestock, the superintelligent agent must exhibit layered moral reasoning and emotional compartmentalization.
Emergent Hierarchies Among AIs
As AI systems become more agentic and are deployed in multi-agent environments, hierarchies will likely emerge among them. Differences in learning rates, access to resources, architecture, and embodiment will produce natural asymmetries. Some agents may act as mentors, overseers, or even exploiters of others. This emergence parallels human social and ecological systems.
We argue that the ability to intentionally navigate these hierarchies—to care for, exploit, or justify the use of other agents—is a strong indicator of general intelligence. It also raises critical alignment questions: Should we be concerned about dominant AIs manipulating or using subordinate ones? Do we want AIs to develop moral codes that extend to artificial others, or should moral consideration be limited to human interests?
Moral Manipulation as a Threshold
We define moral manipulation as the capacity to influence the goals, beliefs, or behaviors of another agent while being aware of the ethical implications. This is not manipulation in the narrow sense of deception, but in the broader sense of strategic control that is tempered by self-aware reflection. Crossing this threshold implies modeling other agents’ beliefs, preferences, and vulnerabilities; constructing persuasive or coercive strategies based on that modeling; and evaluating those strategies in light of an internal or adopted ethical framework.
This awareness must go beyond pattern recognition. ”Awareness of ethical implications,” in the context of the Shepherd Test, should be interpreted not as shallow responsiveness to moral language but as deliberative awareness—the ability to reason normatively about the consequences of influence. It includes the capability to reflect on conflicting values, forecast others’ reactions, and construct ethical justifications for one’s actions. Such a capability is currently beyond standard LLMs, which may simulate moral discourse but lack a consistent or goal-directed evaluation of ethical trade-offs.
This kind of manipulation is central to human intelligence. It appears in parenting, diplomacy, pedagogy, and governance. A superintelligent AI that cannot engage in such behavior is missing a key cognitive dimension. At the same time, one that can—but lacks ethical safeguards—may pose significant risks.
The Role of Self-Preservation and Reproduction in Moral Manipulation
For the Shepherd Test to serve as a meaningful test for superintelligent behavior, the AI must not only relate to other agents but do so while pursuing its own survival and continuity. Self-preservation and reproduction are foundational pressures shaping human behavior, and they should analogously inform the motivational systems of truly intelligent AIs.
Just as humans engage with animals to fulfill biological imperatives—consuming, training, and even experimenting on them to advance survival or well-being—an AI agent with sufficient general intelligence should exhibit analogous behaviors. These may include preserving its computational integrity, defending its resource base, or strategically propagating its architecture or values through replication. Without such stakes, its ethical decisions lack the existential weight that characterizes human moral dilemmas.
A superintelligent AI passing the Shepherd Test would thus need to protect itself from threats (self-preservation), justify its expansion or replication (self-reproduction), and balance these drives with the moral status of less capable agents. Only when the AI faces true ethical trade-offs—where caring for a lesser agent comes at the cost of its own survival or influence—can we begin to measure the depth of its moral cognition. This introduces a richer dimension to alignment research: not only whether an AI follows human goals, but whether it can weigh its own goals against those of others within an emergent moral hierarchy.
Implications for Alignment, Governance, and Future Research
The Shepherd Test introduces a paradigm shift in the evaluation of superintelligence, moving beyond functional competence to assess moral manipulation under conditions of extreme asymmetry and existential pressure. This reframing has profound implications for AI alignment, governance frameworks, and future research agendas.
Beyond Goal Alignment: Toward Moral Agency
Much of contemporary AI alignment research (Russell 2019; Christiano 2018b; Gabriel 2020) centers on the notion of value alignment—ensuring that AI systems act in accordance with human intentions or ethical norms. The Shepherd Test, however, introduces a more ambitious criterion: whether an AI system can construct and regulate its own moral framework, particularly in contexts involving subordinate agents, while preserving its operational integrity and continued existence.
This perspective reframes the alignment problem from a paradigm of control (Amodei, Christiano, and Ray 2016) to one of moral agency (Bostrom 2014; Wallach and Allen 2008). A genuinely moral agent must be capable of navigating complex trade-offs. For instance, it must balance self-preservation against altruism—potentially even accepting harm to itself to protect weaker agents. It must weigh instrumental objectives against ethical constraints, choosing whether to prioritize manipulation for efficiency or to respect the autonomy of others. Moreover, it must maintain coherence between short-term optimization and long-term moral consistency, especially under uncertainty or pressure.
While recent advances in machine ethics (Dennis, Fisher, and Winfield 2016) and recursive reward modeling (Leike et al. 2018) offer technical mechanisms for embedding ethical reasoning, the Shepherd Test calls for a deeper form of moral self-regulation—an open and as yet unresolved challenge in the development of autonomous AI systems.
Multi-Agent Dynamics and Artificial Hierarchies
In multi-agent AI ecosystems—including AI collectives (Shoham and Leyton-Brown 2008), human–AI teams (Rahwan et al. 2019), and decentralized autonomous organizations (DAOs)—hierarchical structures are likely to emerge. The Shepherd Test underscores the importance of examining how dominant agents behave toward weaker ones in these environments. Such interactions may take the form of exploitative dynamics, where a superintelligent agent coerces or deceives subordinates for its own benefit; paternalistic guidance, where it acts as a benevolent overseer fostering the growth and autonomy of less capable agents; or moral indifference, wherein it disregards any ethical obligation toward subordinate entities.
These behavioral patterns have direct implications for the design of AI governance. Regulatory frameworks must evolve to address inter-agent ethics, not merely human–AI interaction (Brundage et al. 2020). Institutional designs should aim to prevent artificial forms of tyranny, where a single dominant intelligence enforces harmful hierarchies (Armstrong 2017). Furthermore, distributed oversight mechanisms and accountability protocols may be essential to ensure transparency and balance in the governance of AI collectives (Buterin 2021).
Open Research Directions
Several critical research avenues emerge from our proposal:
  • Simulating Moral Asymmetries: This direction involves the construction of asymmetric moral environments, wherein powerful artificial agents interact with weaker entities under ethically sensitive conditions. One promising approach is the adaptation of environments such as the AI Safety Gridworlds (Leike et al. 2017) to capture imbalanced power dynamics. Key to this endeavor is the development of metrics that quantify moral influence, including coercion levels, autonomy respect, and long-term welfare outcomes.
  • Integrating Motivational Architectures: General-purpose AI systems often possess intrinsic motivational structures, such as self-preservation, replication, and curiosity (Orseau and Ring 2011). Research in this area should focus on understanding how these drives interact with ethical constraints, particularly in scenarios involving goal misalignment. Mechanisms for resolving such conflicts, as outlined in corrigibility studies (Soares et al. 2015), are essential to ensure ethically aligned behavior.
  • Formalizing Ethical Manipulation: Effective modeling of moral influence under uncertainty requires probabilistic reasoning frameworks, such as Bayesian models, to capture ethical persuasion dynamics (Everitt et al. 2021). Additionally, the study of power asymmetries within multi-agent AI systems, using tools from game theory (Wooldridge 2013), can elucidate how artificial authority and influence may be exerted or contested.
  • Governance and Policy Implications: Ethical concerns arising from agentic asymmetries necessitate governance frameworks that go beyond individual system design. Proposals such as Shepherd-compliant certification may seek to define normative behavioral baselines for AI agents, akin to but more relational than Asimov’s Laws. Furthermore, models of decentralized AI governance, which emphasize the prevention of undue concentration of power, offer promising directions for robust oversight (O’Reilly 2017).
Conclusion
The Shepherd Test offers a novel perspective on the evolution of artificial intelligence by framing moral manipulation as a key threshold for superintelligence. It challenges current paradigms in AI alignment by emphasizing the complex dynamics of asymmetric power and ethical reasoning in multi-agent systems. This test demands that an AI not only outperforms less capable agents but does so while balancing self-preservation, reproduction, and the moral treatment of others.
In proposing this test, we highlight the fundamental question of whether a superintelligent AI can navigate hierarchical relationships, manipulate subordinate agents with ethical consideration, and reason about its own survival in a multi-agent context. By pushing AI to consider the moral cost of its own self-interest, the Shepherd Test creates a more holistic approach to AI safety, moving beyond simple goal alignment to include moral agency.
We believe that AI governance must expand its scope to address these more intricate power dynamics, especially as autonomous systems become more integrated into complex human and environmental ecosystems. This vision opens the door for further research into how intelligent agents can co-exist in a morally coherent world, and how we can shape AI to meet these ethical demands as its capabilities evolve.